OpenStudy (anonymous):

Locate the foci of the ellipse. Show your work. x^2/36+y^2/11=1

OpenStudy (anonymous):

@celinegirl

OpenStudy (anonymous):

if you asked my to guess i would say 6 and squareroot of 11 loll

OpenStudy (anonymous):

no one likes these conic section questions, but they are not that hard

OpenStudy (anonymous):

the center is evidently \((0,0)\) the only thing we need to get is the way it is oriented

OpenStudy (anonymous):



OpenStudy (anonymous):

left?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

because the larger number is under the \(x^2\) terms

OpenStudy (anonymous):

now all we need to know is how many units to the left and right of \((0,0)\) the foci are

OpenStudy (anonymous):

it is \[\sqrt{a^2-b^2}\] which in your case is easy since \(a^2=36,b^2=11\) so it is \[\sqrt{36-11}=\sqrt{25}\]

OpenStudy (anonymous):

aka 5

OpenStudy (anonymous):

what is 5 units to the left of \((0,0)\)?

OpenStudy (anonymous):

(5,0)and(−5,0)